The combustion of octane, C8H18, proceeds according to the reaction

If 5.70 × 102 mol of octane combusts, what volume of carbon dioxide is produced at 25.0 °C and 0.995 atm?
Answer #1

Balanced chemical equation is:

2 C8H18 + 25 O2 ---> 16 CO2 + 18 H2O

From reaction,

mol of CO2 formed = (16/2)*moles of C8H18

= 8*5.70*10^2 mol

= 4560 mol

Given:

P = 0.995 atm

n = 4560 mol

T = 25.0 oC

= (25.0+273) K

= 298 K

use:

P * V = n*R*T

0.995 atm * V = 4560 mol* 0.08206 atm.L/mol.K * 298 K

V = 1.12*10^5 L

Answer: 1.12*10^5 L